# DeepCell: Multiview Representation Learning for Post-Mapping Netlists

### v1

@@ III-E Model pretaining
To train DeepCell, we employ a two-stage training strategy. As illustrated in Fig. 1, in the first stage, we perform random simulation with $15,000$ patterns on PM netlist and record the logic-1 probability $prob_{i},i\in\mathcal{V}^{P}$. Next, we use a simple 3-layer MLP to readout the functional embeddings as the predicted probability. The training loss for this stage is defined in Eq. (8).
$\begin{split}\hat{prob}_{i}&=MLP(hf_{i}),~{}i\in\mathcal{V}^{P}\\ loss_{prob}&=L1Loss(prob_{i},prob_{i}^{{}^{\prime}})\\ \end{split}$ (8)
In the second stage, we refine the cell embeddings by incorporating representations from the AIG view. We utilize a pretained and frozen AIG encoder to provide rich contextual information from the AIG. The training objective in this stage is to reconstruct the functionality of masked cells using information from their neighboring cells and the global AIG perspective. Accordingly, we define the MCM training loss in Eq. (9), where the model is trained to recover the functional embeddings of the masked cells.
$loss_{mcm}=L1Loss(H_{i}^{P},H_{i}^{P^{\prime}})$ (9)
@@
$\begin{split}L_{stage1}&=loss_{prob}\\ L_{stage2}&=w_{prob}\cdot loss_{prob}+w_{mcm}\cdot loss_{mcm}\\ \end{split}$ (10)
@@ IV Experiments
In this section, we train our DeepCell to learn the general representations of standard cells in the post-mapping netlists.
@@ IV-A1 Data Preparation
We collect the open-source verilog designs from GitHub and the open-source circuit benchmarks [14, 15, 16]. All circuits are designed with Skywater 130nm as the target process library [17], using commercial logic synthesis tool Design Compiler command compile_ultra to obtain the corresponding PM netlist. To accelerate model pretraining, we randomly extract the sub-circuits within 5,000 cells. Then, we use ABC strash command to obtain the corresponding AIGs. Consequently, there are totally 83,155 PM netlists and corresponding AIGs. We split the dataset into 80% for training and 20% for testing.
@@ IV-A2 Evaluation Metrics
Additionally, we assess model performance on mask circuit modeling tasks, where the average reconstruction error (RE) is defined as Eq. (12), where all the select nodes $p$ are in set $\mathcal{S}$ and the masked hop is $\mathcal{M}(p)$.
@@ IV-A3 Model Implementation and Training
The Transformer model used to refine cell embeddings consists of 4 Transformer blocks, each with 8 attention heads. After encoding, each cell is represented by a 128-dimensional structural embedding and a 128-dimensional functional embedding. DeepCell is pretrained for 60 epochs in Stage 1 and an additional 60 epochs in Stage 2 to ensure convergence. The pretraining process is conducted with a batch size of 128 using 8 Nvidia A800 GPUs. We employ the Adam optimizer with a learning rate of $10^{-4}$.
@@ IV-B Experimental Results
IV-B1 Effect of Mask Modeling
We investigate the optimal settings for Mask Circuit Modeling (MCM) by exploring its two key hyperparameters: the number of selected nodes to be masked, $|\mathcal{S}|$, and the size of the masked hop, $k$. To evaluate performance, we pretrain the model using various hyperparameter combinations, where $\mathcal{|S|}=\theta\cdot|\mathcal{V}^{P}|$ and $\theta=1\%,5\%,10\%,20\%$ of total nodes in PM netlist, with hop sizes of $k=4$ or $k=6$.
TABLE I: Effect of Mask Modeling
@@ IV-B1 Effect of Mask Modeling
Table I presents the results for different values of $\theta$ and $k$. First, the reconstruction error (RE) increases with a larger masking hop size. For example, when $\theta=0.05$, the RE for $k=6$ is 0.0257, which is 164.95% higher than that for $k=4$ (RE=0.0097). Second, masking a smaller number of nodes consistently reduces both the RE and PE. However, using a smaller $\theta$ makes the task less challenging and diminishes its effectiveness as a pretraining objective. Based on these observation, we select $\theta=0.05$ and $k=4$ as a trade-off between task complexity and model performance in the following experiments.
IV-B2 Effect of Multiview Learning
@@
We investigate the impact of incorporating the AIG view on training our PM netlist representation model. To evaluate the representation capability of the DeepCell framework, we use the PE as the primary metric. Specifically, we compare the full multiview MCM training strategy (w/ multiview) with a baseline that uses only the PM encoder without refining embeddings through multiview training (w/o multiview). The RE metric is not available in the w/o multiview setting. It is important to note that the DeepCell w/o multiview setting reflects the representation capability of DeepGate2 [2] on PM netlists, since our DeepCell utilizes a similar aggregator in the PM encoder.
To the best of our knowledge, no prior work has focused on learning a general representation of PM netlists composed of standard cells. In addition to evaluating DeepCell, we also employ general Graph Neural Network (GNN) models, such as GCN [12] and GAT [13], as PM encoders to investigate the impact of MCM training.
TABLE II: Effect of Multiview Learning
PE RE
GCN [12] w/o multiview 0.0956 -
w/ multiview 0.0529 0.0571
GAT [13] w/o multiview 0.1466 -
w/ multiview 0.0682 0.0869
DeepCell w/o multiview 0.0564 -
w/ multiview 0.0323 0.0097
Table II compares the performance of the models with and without multiview training. First, compared to other PM encoders, whether using multiview training or not, DeepCell outperforms alternatives such as GCN [12] and GAT [13], achieving the lowest PE of 0.0323. Second, DeepCell effectively reconstructs functional embeddings by leveraging surrounding cell embeddings and additional gate embeddings from the AIG, achieving an average RE of 0.0097. Third, all models with multiview training outperform their counterparts without multiview training in terms of PE. For example, DeepCell with multiview training reduces PE by 42.7%, from 0.0564 to 0.0323. These results demonstrate the effectiveness of the proposed multiview representation learning mechanism for PM netlists.
@@ V Downstream Task: Functional ECO
In this section, we combine our DeepCell with the open-source EDA tools and apply our model to practical EDA tasks: functional ECO. Our DeepCell provides probability for nodes in the original circuit and provides guidance for ECO to find candidate patch signals. This method effectively leaves higher quality and more likely to be a patch signal, making the ECO solution more rapid and higher quality.
@@ V-A Preliminary
In order to clarify the role of the model in the algorithm, we first briefly describe the method in [19]. In the original SAT-based solution, for the POs reachable from the target, the method simultaneously identifies the PIs contained in these POs in the original netlist and the target netlist. Then find all signals in the original circuit, which are not in the transitive fanout cone of the targets and whose support is contained in the calculated set of PIs. These candidate signals are sorted by cost in ascending order and a fixed number of nodes (default is the top 5,000) are selected. Then used the SAT-based solution for the single-target ECO problem proposed in [20] to prove whether there is a solution.
After adding a large number of candidate signals, it is very expensive to use the positive and negative remainders of target to construct the miter circuit and solve it, which is also the reason why the number of candidate signals cannot be selected in large quantities. It is necessary to reduce the number of candidate signals at the outset.
@@ V-B Model Finetuning
DeepCell is pretrained to capture the correlation between masked PM netlist and complete AIG. Intuitively, the original circuit that requires patch insertion in ECO task is treated as PM netlist with masking, while the golden circuit provide the complete view. We then finetune DeepCell to effectively learn and represent the relationship between these two circuits.
Figure 4: Select candidate signal with DeepCell
@@
Therefore, we finetune DeepCell to reduce the space of candidate signals in the ECO process. Specifically, given a PM netlist $\mathcal{G}^{P}$, we randomly select a cell $p$ within the netlist and remove its entire driven cone from this node to PI. The removed area, $\mathcal{M}(p)$, is treated as the ground-truth path in ECO. Consequently, the patch-removed netlist is denoted as $\widetilde{\mathcal{G}}^{P}$, while the corresponding AIG netlist for the original PM netlist is denoted as $\mathcal{G}^{A}$. Next, we use the PM encoder to encode both patch-removed netlist $\widetilde{\mathcal{G}}^{P}$ and the original netlist $\mathcal{G}^{P}$, as denoted in Eq. (13). The AIG encode is employed to encode the AIG netlist $\mathcal{G}^{A}$, which serves as the reference ground-truth implementation. The embeddings of the patch-removed netlist are then refined using the AIG embeddings, as described in Eq.(7), to produce the refined embeddings $\widetilde{\mathbf{H}}^{P}$.
$\begin{split}\mathbf{H}^{A}&=DeepGate2(\mathcal{G}^{A})\\ \mathbf{H}^{P}&=encoder(\mathcal{G}^{P})\\ \widetilde{\mathbf{H}}^{P}&=encoder(\widetilde{\mathcal{G}}^{P})\\ \{\mathbf{H}^{A^{\prime}},\widetilde{\mathbf{H}}^{P^{\prime}}\}&=\mathcal{T}(\{\mathbf{H}^{A},\widetilde{\mathbf{H}}^{P}\})\end{split}$ (13)
TABLE III: Performance Comparison between w/ DeepCell and w/o DeepCell on ICCAD’17 CAD Contest Benchmarks
Circuit name Circuit information w/o DeepCell w/ DeepCell
PI PO gate(F) gate(G) tatget cost gate time(s) cost gate time(s)
unit 1 3 2 6 6 1 4 1 0.37 4 1 0.37
unit 2 157 64 1120 1219 1 17 4 0.92 17 4 1.01
unit 3 411 128 2074 1929 1 80 3 0.46 80 3 0.4
unit 4 11 6 75 77 1 32 5 0.49 32 2 0.55
unit 5 450 282 24357 21056 2 47 30 46.3 47 29 15.39
unit 6 99 128 13828 11812 2 - - - - - -
unit 7 207 24 2944 1721 1 284 2 8.5 284 2 6.88
unit 8 179 64 2513 3337 1 78 4 4.1 78 3 3.17
unit 9 256 245 5849 4657 4 50 35 1.33 50 26 1.7
unit 10 32 129 1581 1956 2 - - - - - -
unit 11 48 50 2057 2160 8 4142 1063 7200 2312 746 6975.3
unit 12 46 27 13804 821 1 104 1 1.02 104 1 0.71
unit 13 25 39 369 426 1 3467 9 1.5 3467 9 1.33
unit 14 17 15 1981 1006 12 95 41 5.43 95 41 6.72
unit 15 198 14 1886 2262 1 191 11 1.93 191 11 0.73
unit 16 417 214 2371 9324 2 278 15 16.58 204 9 14.67
unit 17 136 31 2910 2052 8 434 79 5.04 434 63 6.01
unit 18 245 100 4860 3881 1 18 1 5.14 18 1 2.44
unit 19 99 128 13349 10787 4 - - - - - -
unit 20 1874 7105 30876 34002 4 136 6 0.6 120 5 1.12
Geomean 112.36 8.96 3.72 105.83 7.60 3.19
Imp. 1 1 1x 0.06 0.15 0.85x
The finetuning tasks are defined as patch reconstruction and driven signal identification. First, in the ideal scenario, the finetuned model should recover the embeddings of the patch with the correct functionality, as informed by the AIG view. We treat the embeddings of the original netlist (prior to patch removal) as the ground truth and define the patch reconstruction loss, $loss_{pr}$ as specified in Eq. (14) (see Fig. 3). We finetune DeepCell using only 8,000 circuits over 10 epochs with loss function $L=loss_{pr}+loss_{drv}$.
$loss_{pr}=L1Loss(H_{i}^{P^{\prime}},\widetilde{H}_{i}^{P^{\prime}}),~{}i\in\mathcal{M}(p)$ (14)
Second, in the driven signal identification task, we utilize the refined embeddings to predict whether a given cell is driven by the selected cell $p$.
$loss_{drv}=BCELoss(MLP(\{\widetilde{H}_{i}^{P^{\prime}},\widetilde{H}_{p}^{P^{\prime}}\}),0/1),i\in\mathcal{V}^{P}$ (15)
As shown in Fig. 4, with DeepCell intervention, according to the obtained probability, we remove signals that are almost impossible to drive as targets, so that the number of candidate signals is greatly reduced, so that the situation that needs to be considered in the subsequent steps is controlled in a small range. This operation greatly reduces the running time.

### v2

@@ III-E Model pretaining
To effectively learn rich multiview embeddings, we employ a two-stage training strategy. In the first stage, the GNN-based encoders are pretrained by predicting the logic-1 probability under random simulation, which provides rich supervision containing both structural and semantic information of Boolean logic [3]. Specifically, we perform random simulation with $15,000$ patterns on PM netlist and record the logic-1 probability $prob_{i},i\in\mathcal{V}^{P}$. Next, we use a simple 3-layer MLP to readout the embedding of each cell in PM netlist as the predicted probability. The training loss of PM encoder is defined in Eq. (8).
$\begin{split}\hat{prob^{P}_{i}}&=MLP(hf_{i}),~{}i\in\mathcal{V}^{P}\\ L_{prob}^{P}&=L1Loss(prob_{i},\hat{prob^{P}_{i}})\\ \end{split}$ (8)
We train the AIG encoder in a similar manner. The training loss of AIG encoder is defined in Eq. (9).
$\begin{split}\hat{prob^{A}_{i}}&=MLP(hf_{i}),~{}i\in\mathcal{V}^{P}\\ L_{prob}^{A}&=L1Loss(prob_{i},\hat{prob^{A}_{i}})\\ \end{split}$ (9)
In the second stage, we refine the cell embeddings by incorporating representations from the AIG view. We utilize the pretained and frozen AIG encoder to provide rich contextual information from the AIG. The training objective in this stage is to reconstruct the functionality of masked cells using information from their neighboring cells and the global AIG perspective. Accordingly, we define the MCM training loss in Eq. (10), where the model is trained to recover the functional embeddings of the masked cells.
$L_{mcm}=L1Loss(H_{i}^{P},H_{i}^{P^{\prime}})$ (10)
Alternatively, if DeepCell is employed to refine AIG embeddings and the PM encoder $\Phi^{P}$ is frozen. The loss function is modified to:
$L_{mcm}=L1Loss(H_{i}^{A},H_{i}^{A^{\prime}})$ (11)
@@
$\begin{split}L_{stage1}&=L_{prob}^{P}+L_{prob}^{A}\\ L_{stage2}&=w_{prob}\cdot(L_{prob}^{P}+L_{prob}^{A})+w_{mcm}\cdot L_{mcm}\\ \end{split}$ (12)
@@ IV-A1 Data Preparation
We employ ForgeEDA [32] and DeepCircuitX [2] datasets, which collect open-source repository-level designs from GitHub. We randomly extract sub-circuits (both PM netlists and AIGs) containing up to 4,096 nodes for efficient model pre-training. It is important to note that DeepCell trained on small-scale sub-circuits can be generalized to larger circuits with acceptable runtime, as demonstrated in the practical downstream tasks discussed in Sec. V and Sec. VI.
For model training and validation, we utilize three open-source technology libraries (Skywater130nm [33], GSCL45nm [34], and GlobalFoundries(GF)180nm [35]) to generate post-mapping netlists using commercial tool with a similar distribution (ID). Additionally, we prepare the testing dataset using another technology library, Nangate45nm [36], which serves as an out-of-distribution (OOD) set. The statistics of our datasets are provided in TABLE I, where Dist. indicates whether the set has the same distribution as the training set.
TABLE I: The statistics of DeepCell datasets
Dataset Dist. Libraries # Circuits
Training ID Skywater130nm, GSCL45nm, GF180nm 16,072
Validation ID Skywater130nm, GSCL45nm, GF180nm 4,019
Testing OOD Nangate45nm 1,828
@@ IV-A2 Evaluation Metrics
Additionally, we assess model performance on mask circuit modeling tasks, where the average reconstruction error (RE) is defined as Eq. (14), where all the select nodes $p$ are in set $\mathcal{S}$ and the masked hop is $\mathcal{M}(p)$. Lower values ($\downarrow$) indicate better performance for both metrics.
@@ IV-A3 Model Implementation and Training
The Transformer-based model used to refine cell embeddings consists of 4 Transformer blocks, each with 8 attention heads. After encoding, each cell is represented by a 128-dimensional structural embedding and a 128-dimensional functional embedding. DeepCell is pretrained for 60 epochs in Stage 1 and an additional 60 epochs in Stage 2 to ensure convergence. The pretraining process is conducted with a batch size of 128 using 8 Nvidia A800 GPUs. We employ the Adam optimizer with a learning rate of $10^{-4}$.
@@ IV-B Experimental Results
IV-B1 Hyperparameter Analysis
We investigate the optimal settings for model pretraining by exploring two key hyperparameters in MCM: the number of selected nodes to be masked, $|\mathcal{S}|$, and the size of the masked hop, $k$. To evaluate performance, we pretrain the model using various hyperparameter combinations, where $\mathcal{|S|}=\theta\cdot|\mathcal{V}^{P}|$ and $\theta=1\%,5\%,10\%,20\%$ of total nodes in PM netlist, with hop sizes of $k=4$ or $k=6$.
TABLE II: Effect of Different Hyperparameter
@@ IV-B1 Hyperparameter Analysis
TABLE II presents the results for different values of $\theta$ and $k$. First, the reconstruction error (RE) increases with a larger masking hop size. For example, when $\theta=0.05$, the RE for $k=6$ is 0.0257, which is 164.95% higher than that for $k=4$ (RE=0.0097). Second, masking a smaller number of nodes consistently reduces both the RE and PE. However, using a smaller $\theta$ makes the task less challenging and diminishes its effectiveness as a pretraining objective. Based on these observation, we select $\theta=0.05$ and $k=4$ as a trade-off between task complexity and model performance in the following experiments.
IV-B2 Effect of Multiview Learning on PM Netlist
@@
We investigate the impact of incorporating the multiview information on PM netlist representation learning. Specifically, we compare the full multiview MCM training strategy (w/ multiview) with a baseline that uses only the PM encoder without refining embeddings through multiview training (w/o multiview). The RE metric is not available in the w/o multiview setting. To the best of our knowledge, no prior work has focused on learning general-purpose representations of post-mapping (PM) netlists composed of standard cells. To evaluate the effectiveness and generalization ability of our proposed Mask Circuit Modeling (MCM) training strategy, we incorporate different GNN-based aggregators: Convolutional Sum (Conv. Sum) [37], Attention [38] and the aggregator in DeepGate2 (DG2) [5], into PM encoders within the DeepCell framework.
TABLE III: Effect of Multiview Learning on PM Netlist Encoder
Aggr Multview Dist. PE Red. RE
Conv. Sum w/o ID 0.0344 - -
w/ ID 0.0312 9.41% 0.0081
w/o OOD 0.0380 - -
w/ OOD 0.0360 5.29% 0.0096
Attention w/o ID 0.0337 - -
w/ ID 0.0318 5.65% 0.0076
w/o OOD 0.0350 - -
w/ OOD 0.0325 7.04% 0.0087
DG2 w/o ID 0.0314 - -
w/ ID 0.0286 8.92% 0.0082
w/o OOD 0.0316 - -
w/ OOD 0.0295 6.68% 0.0089
TABLE III compares the performance of the models with (w/) and without (w/o) multiview training. The column ‘Red.’ shows the PE reduction achieved by incorporating multiview information. First, compared to other PM encoders, whether using multiview training or not, the DG2 aggregator outperforms alternatives with lowest PE of 0.0286. We employ such aggregator in the following applications. Second, all models with multiview training outperform their counterparts without multiview training in terms of PE. For example, Conv. Sum w/ multiview reduces PE by 9.41%, from 0.0344 to 0.0312. Third, our PM netlist encoder performs similarly on both ID and OOD datasets, highlighting the generalization ability of DeepCell across different technology libraries.
IV-B3 Effect of Multiview Learning on AIG
We further explore the impact of multiview learning on enhancing AIG representations. Specifically, we compare the performance of AIG encoders with (w/) and without (w/o) multiview training using the PE metric. Four representative AIG encoders—PolarGate [10], HOGA [7], DeepGate2 [5], and DeepGate3 [8]—along with a general GCN model [31], are evaluated in the following experiments. As shown in TABLE IV, incorporating multiview training significantly improves model performance. For example, DeepGate2 achieves a PE of 0.0346 without multiview training, but with multiview training, the PE is reduced by 28.03%, reaching 0.0249. This demonstrates the considerable benefit of adding information from the PM view.
Furthermore, since the outputs of the GNN-based AIG encoder are treated as AIG embeddings within the DeepCell framework (see Fig. 1), the runtime and memory usage for AIG representation learning during inference remain identical to those of the baseline models. This ensures that the introduction of multiview learning does not incur any additional computational overhead, while still delivering substantial performance improvements.
TABLE IV: Effect of Multiview Learning on AIG Encoder
Model Multview PE Red. RE
GCN [31] w/o 0.0599 - -
w/ 0.0540 9.85% 0.0081
PolarGate [10] w/o 0.0318 - -
w/ 0.0309 2.83% 0.0076
HOGA [7] w/o 0.0704 - -
w/ 0.0685 2.70% 0.0084
DeepGate2 [5] w/o 0.0346 - -
w/ 0.0249 28.03% 0.0087
DeepGate3 [11] w/o 0.0263 - -
w/ 0.0244 7.22% 0.0093
@@ V-A Preliminary
Functional Engineering Change Orders (ECO) is an essential task in EDA, where an existing circuit design is updated to meet new target specifications without requiring a full redesign. Specifically, given an original PM netlist $\widetilde{\mathcal{G}}^{P}$ and the target patch location, functional ECO focuses on adding a patch to construct another netlist $\mathcal{G}^{P}$, which meets the functionality of the golden design. The method introduced in [16] for functional ECO wins the ICCAD contest in this domain. The algorithm first identifies the primary outputs (POs) that are reachable from the target patch location. For each reachable PO, the method then calculates the complete driven cones, identifying the candidate signals that could potentially be involved in the patch generation. Finally, the selected candidate signals are used to prove whether there is a solution using SAT-based approach [39].
However, the computational cost of generating the patch and verifying the solution using SAT solvers becomes prohibitively expensive as the number of candidate signals increases. Although [16] restricts the SAT-based approach to generate the patch using only a fixed number of low-cost candidate signals in order to save runtime, this constrained set often leads to situations where certain ECO problems cannot be solved. The challenge, therefore, lies in efficiently reducing the number of candidate signals considered, while still ensuring that the ECO problem can be solved within acceptable computational constraints.
@@ V-B Model Finetuning
DeepCell is employed to effectively narrow down the candidate signal space by learning which signals are most likely to be relevant for the patch generation task. Intuitively, the original circuit that requires patch insertion in ECO task is treated as PM netlist with masking, while the golden circuit provide the complete view. Therefore, we finetune DeepCell to predict the driven signals of the patch with the view from both original PM netlist and golden AIG.
@@
As shown in Fig. 3, given a PM netlist $\mathcal{G}^{P}$, we randomly select a cell $p$ within the netlist and remove its entire driven cone from this node to PI. The removed area, $\mathcal{M}(p)$, is treated as the ground-truth path in ECO. Consequently, the patch-removed netlist is denoted as $\widetilde{\mathcal{G}}^{P}$, while the corresponding AIG for the PM netlist $\mathcal{G}^{P}$ is denoted as $\mathcal{G}^{A}$. Next, we use the PM encoder $\Phi^{P}$ to encode both $\widetilde{\mathcal{G}}^{P}$ and $\mathcal{G}^{P}$, obtaining the corresponding embeddings $\widetilde{\mathbf{H}}^{P}$ and $\mathbf{H}^{P}$. The AIG encode $\Phi^{A}$ is employed to encode the AIG netlist $\mathcal{G}^{A}$, where the AIG embeddings $\mathbf{H}^{A}$ serves as the view from golden implementation. The embeddings of the patch-removed PM netlist are then refined using the AIG embeddings using Transformer Blocks to produce the refined tokens (see Eq. (15)).
$\begin{split}\{\mathbf{H}^{A^{\prime}},\widetilde{\mathbf{H}}^{P^{\prime}}\}&=\mathcal{T}(\{\mathbf{H}^{A},\widetilde{\mathbf{H}}^{P}\})\end{split}$ (15)
Figure 3: DeepCell finetuning for functional ECO
The finetuning tasks are defined as patch reconstruction (Reconstruction) and driven signal prediction (Driven Pred.). In the first task, we treat the embeddings of the golden PM netlist $\mathcal{G}^{P}$ as the ground truth and define the reconstruction loss $L_{rec}$. The goal is to minimize the difference between the reconstructed patch-removed PM netlist and the original golden PM netlist. In the second task, we focus on driven signal prediction. Here, the refined embeddings $\widetilde{\mathbf{H}}^{P^{\prime}}$ are used to predict whether a certain cell in the netlist is driven by the cell $p$ at the patch insertion location.
$\begin{split}L_{rec}&=L1Loss(H_{i}^{P^{\prime}},\widetilde{H}_{i}^{P^{\prime}}),~{}i\in\mathcal{M}(p)\\ L_{drv}&=BCE(MLP(\{\widetilde{H}_{i}^{P^{\prime}},\widetilde{H}_{p}^{P^{\prime}}\}),0/1),i\in\mathcal{V}^{P}\end{split}$ (16)
We finetune the pretrained DeepCell using 8,000 circuits only over 10 epochs (about 2 GPU hours) with loss function $L=L_{rec}+L_{drv}$. To ensure efficiency during inference while accepting a slight trade-off in accuracy, we implement the Transformer Blocks using Linear Transformer [40]. During inference, we only process the original PM netlist $\widetilde{\mathcal{G}}^{P}$ (the patch-removed netlist used during finetuning) and the golden AIG $\mathcal{G}^{A}$. Our model produces a list of floating-point numbers, each representing the probability that a given cell should be included in the candidate signal set.